OpenStudy (anonymous):

what is the discontinuity of the function f(x)= x^2-4x-12/x+2 ? What is the discontinuity and zero of the function f(x)= 3x^2+x-4/x-1

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

If you need the zeros of the function, \[f(x)=\frac{ x^2-4x-12}{x+2}\] You must solve, \[f(x)=0\Rightarrow x^2-4x-12=0\] For the second question, you must solve two things: - The zeros, \[3x^2+x-4=0\] - The point of discontinuity, \[x-1=0\Rightarrow x=1\] Can you complete the problem?

OpenStudy (john_es):

Remember, that any rational function of the form, \[\frac{P(x)}{Q(x)}\] it's not defined when \[Q(x)=0\] So those points x where Q(x)=0, are disconuity points.

OpenStudy (anonymous):

woops I made a mistake for the first one I need the discontinuity not the zeroes.

OpenStudy (anonymous):

here are the questions and choices.

OpenStudy (john_es):

For the first question if you need the discontinuity point, as you see, it is easy to find, \[x+2=0\Rightarrow x=-2\] You should choose (-2,8), but I don't understand why they add an 8 to the answer.

OpenStudy (anonymous):

because we are supposed to use Geogebra, a graphing tool, to help us do this.

OpenStudy (john_es):

For the second question, the last option is the correct one.

OpenStudy (john_es):

So, you should use geogebra to check, at least, these solutions.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

the first one is right

OpenStudy (anonymous):

and so is the 2nd one